INFRARED DIODE SPECTROSCOPIC INVESTIGATION OF THE FOUR C-H STRETCHING VIBRA-TIONAL MODES OF PROPYLENE OXIDE

\underline{CHEOLHWA KANG}, ZHENG SU, and YUNJIE XU, Department of Chemistry, University of Alberta, Edmonton, AB T6G 2G2 Canada.

Propylene oxide is a simple chiral building block and is a prototype for theoretical and experimental studies of chiral recognition effects. The low resolution infrared vibrational spectra in the C-H stretching region of propylene oxide in the gas phase^{a} and in a free jet expansion^{b} were reported previously. In this talk, we will present the first high resolution infrared spectra of four C-H stretching vibrational modes. The spectra have been measured using a rapid scan mid-infrared lead salt diode laser spectrometer equipped with an astigmatic multi-path sample cell^{c}. Analyses of the observed spectra show that there are severe perturbations associated with the first vibrationally excited states. The detailed analyses will be discussed.

aF. Winther and D. O. Hummel,Spectro. Acta, 1969, 25A, 417.

bN. Borho and M. Suhm, unpublished results.

cZ. Su, W. Tam, and Y. Xu, J. Chem. Phys., 2006, 124, 024311/1-9.